OpenStudy (anonymous):

1/12x - 2/3= 1/2(3/4- 1/3x)

OpenStudy (anonymous):

is this the one where you have to make all the denominators common?

OpenStudy (anonymous):

i dont understand how to do this

OpenStudy (zepp):

http://puu.sh/ro1z There, you have to distribute the 1/2

OpenStudy (campbell_st):

is this the problem...? \[\frac{1}{12} x -\frac{2}{3} = \frac{1}{2} (\frac{3}{4} - \frac{1}{3} x)\]

OpenStudy (anonymous):

yes

OpenStudy (zepp):

http://puu.sh/ro2o Then you can start moving thing around, term with a x together and constants together

OpenStudy (anonymous):

and then you ultiply 1/2 on the other side correct?

OpenStudy (zepp):

You have to develop it first, as shown in the picture I posted above.

OpenStudy (campbell_st):

ok... multiply both sides by 2 to get rid of 1/2 on the Right \[\frac{1}{6} x - \frac{4}{3} = \frac{3}{4} - \frac{1}{3} x\] add 1/3 x to both sides \[(\frac{1}{3} + \frac{1}{6}) x - \frac{4}{3} = \frac{3}{4}\] add 4/3 to both sides \[\frac{1}{2} x = \frac{3}{4} + \frac{4}{3}\] \[\frac[1}{2}x = \frac{25}{12}\] multiply both sides by 2 \[x = \frac{25}{6}\] or \[x = 4\frac{1}{6}\]